Research Scenario: By using words per sentence and characters per words, we can create formulas for the Flesch-Kincaid gradeQuestion: Use the prediction equation from the regression model for the following: Given a page with an average words per senQuestion: When analyzing the correlation between characters per word (x-variable) and the F-K Grade Level (y-variable), the sQuestion: Analyze the correlation between words per sentence and characters per word. O The correlation is weak and positive.Question: Analyze the correlation between words per sentence and the F-K Grade Level. The correlation is weak and positive. O
Answer #1

We input the given data set in R programming software and use the lm(.), cor(.) and summary(.) functions to answer the questions given here.

(i) The R-squared value is 0.10. Hence, 10% of the variance in characters per word is accounted for by words per sentence.
(ii) Here, x = Words/Sentence, y = F-K Grade level. The regression equation is: \hat{y} = 0.680 + (0.334) x . Now, for x = 10, predicted value of y = 0.680 + (0.334 * 10) = 4.022 = 4.
(iii) Now, x = Characters/Word, y = F-K Grade level. The slope in the regression equation is = 4.305. Hence, answer is True.
(iv) Correlation between Words/Sentence and Characters/Word = 0.312. Hence, the correlation is moderate and positive.
(v) Correlation between Words/Sentence and F-K Grade level = 0.905. Hence, the correlation is strong and positive.